Suppose the production function for a firm is given by: q=4L0.5K0.25. If the firm currently has 10 units of capital (K) and 10 units of labor (L), then calculate the Marginal Rate of Technical Substitution (MRTSLK).

(Round to the nearest two decimal places if necessary.)
Answer #1

Here; q = 420.560.25 on q= 4 1/2 kih Pepe = 2(a) Where c an is the partial differentiation of OF & with respect to t MPL =Similarly au detik Ape = 44/ 28 (k 1/4) ak 2 427 = 11 1 Trily MPR = 12 1314 Now PARTS = Tape Taple - 2 klly K 34 x 2 (1/2 - 2